The sum of 3 consecutive multiples of is 333. Find the multiples

Answer by stanbon(75887)

The sum of 3 consecutive multiples of is 333. Find the multiples
---------
1st: 3x
2nd: 3(x+1)
3rd: 3(x+2)
---------------
3(x+x+1+x+2) = 333
3x+3 = 111
3x = 108
x = 36
----
3x = 108
3(x+1) = 111
3(x+2) = 114
-------------------
Cheers,
Stan H.
----------